OpenStudy (anonymous):

Let f(x) =\frac{2x^2+2 x -40}{x-4} Show that f(x) has a removable discontinuity at x=4 and determine what value for f(4) would make f(x) continuous at x=4.

OpenStudy (kainui):

So if you can factor the top part out and remove an x-4 term from the top, you'll have the exact same graph except with the hole filled in. Plug in 4, and you'll see exactly where it should be.

OpenStudy (gorv):

2x^2+2 x -40=2x2+10x-8x-40 2x(x+5)-8(x+5) (x+5)(2x-8) (x+5)*2*(x-4) now denominator is (x-4) so cancel out x-4 =2(x+5) at x=4 =2*(4+5)=2*9=18 @abjectbrownE

OpenStudy (anonymous):

thanks man

OpenStudy (nincompoop):

you can make the equation continuous so long as your denominator will not be zero you can achieve this by factoring the numerator so that a part of it matches your denominator